OpenStudy (anonymous):

Can someone please explain completing the square? The question is: "Toni is solving this equation by completing the square" ax^2+bx+c=0 (where "a" is greater than or equal to 0) Step 1: ax^2+bx=-c Step 2: x^2+b/a x=-c/a Step 3: ? (I have to figure out what the third step is)

OpenStudy (wolf1728):

Here's how it is done: To complete the square: Move the "non X" term to the right: ax^2+bx+c=0 ax² + bx = -c Divide the equation by the coefficient of X² which in this case is 'a' x² + bx/a = -c/a Now for "completing the square" we: take the coefficient of x which is b/a divide that by 2 b/2a square that number b²/4a² then add it to both sides of the equation x² + bx/a +b²/4a² = -c/a +b²/4a² Take the square root of BOTH sides of the equation

OpenStudy (wolf1728):

It's a lot easier to see when using real numbers: x² + 12x - 8=0 x² + 12x = 8 take the coefficient of X which is 12 divide it by 2 then square that number: 12/2 = 6 and 6² = 36 add 36 to both sides of the equation x² + 12x +36 = 8 +36 Take the square root of BOTH sides of the equation (x+6) = sqrt(44) x = sqrt(44) -6 x = 0.6332496 x = -12.6332496 Yes, it's just that simple. LOL

OpenStudy (anonymous):

Thanks! I appreciate it!